# mystyle backgroundcolor=white!10

### v1

@@ IV Finite-time Reach-avoid Verification
Definition 4
The switched stochastic discrete-time system, which is built upon system (1), is a quadruple (^$(\widehat{\mathcal{L}},\widehat{\mathcal{X}},\bm{x}_{0},\widehat{\bm{f}})$ with the following components:
@@
    ^=$\widehat{\mathcal{L}}=\{1,2,3\}$ is a set of three locations;
@@
    with $\widehat{\bm{f}}_{1}(\bm{x},\bm{\theta})=\bm{f}(\bm{x},\bm{\theta})$, $\widehat{\bm{f}}_{2}(\bm{x},\bm{\theta})=\bm{x}$ and $\widehat{\bm{f}}_{3}(\bm{x},\bm{\theta})=\bm{x}$, and $1_{\widehat{\mathcal{X}}_{i}}(\bm{x})$ is the indicator function of the set $\widehat{\mathcal{X}}_{i}$, i.e., $11_{\widehat{\mathcal{X}}_{i}}(\bm{x})=$ 1 if $\bm{x}\in\widehat{\mathcal{X}}_{i}$; otherwise, $01_{\widehat{\mathcal{X}}_{i}}(\bm{x})=$ 0,
@@
    $\widehat{\Omega}=\{\bm{x}\in\mathbb{R}^{n}\mid\bm{x}=\bm{f}(\bm{x}_{0},\bm{\theta}),\bm{x}_{0}\in\mathcal{X},\bm{\theta}\in\Theta\}\cup\mathcal{X};$
@@
$\begin{split}&\bm{x}(l+1)=\widehat{\bm{f}}(\bm{x}(l),\bm{\theta}(l)),\forall l\in\mathbb{N},\\ &\bm{x}(0)=\bm{x}_{0}\in\widehat{\mathcal{X}}.\end{split}$ (8)
The trajectory of system (8), induced by initial state $\bm{x}_{0}\in\widehat{\mathcal{X}}$ and disturbance policy $\pi\in\Omega$, is denoted by $\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$.
@@
IV-A Sufficient Conditions for Upper Bounds
@@
Theorem 3
@@
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathcal{X}_{r}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}_{r}\end{cases},$ (9)
@@
$\begin{cases}v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\wedge\beta\in[0,1]\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha\in(0,1)\wedge\beta\in[0,1]\end{cases}.$
Proof:
@@
$\begin{cases}\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\widehat{\mathcal{X}}\\ v(\bm{x})\geq 1_{\mathcal{X}_{r}}(\bm{x})&\forall\bm{x}\in\widehat{\mathcal{X}}\end{cases}.$ (10)
@@
Thus, $\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}=\mathbb{E}^{\infty}[1_{\mathcal{X}_{r}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))]\leq\alpha^{-N}v(\bm{x}_{0})+\alpha\beta\frac{(1-\alpha^{-N})}{\alpha-1}.$ According to Lemma 2, we have the conclusion.
@@
On the other hand, the analysis for the upper bounds in Theorem 3 is similar to the one in Theorem 1. When $\alpha=1$ and $\beta=0$, the upper bound is $v(\bm{x}_{0})$, which is an upper bound of the probability $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}$. Besides, since $v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}=\alpha^{-N}(v(\bm{x}_{0})-\frac{\alpha\beta}{\alpha-1})+\frac{\alpha\beta}{\alpha-1}$, we conclude that when $\alpha\in(0,1)$ and $\beta\in[0,1]$, if $v(\bm{x}_{0})=\frac{\alpha\beta}{\alpha-1}$, $v(\bm{x}_{0})=\frac{\alpha\beta}{\alpha-1}$ is also an upper bound of the probability of reaching the target set $\mathcal{X}_{r}$ eventually while staying within the safe set $\mathcal{X}$ before the first target hitting time; if $v(\bm{x}_{0})<\frac{\alpha\beta}{\alpha-1}$, $\lim_{N\rightarrow\infty}v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}=-\infty$, which implies $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}=0$ and thus $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}=0$.
Remark 3
In Theorem 3, $\alpha$ is limited to be in $(0,1]$. We can obtain upper bounds for the case with $\alpha>1$ according to Theorem 3 in Chapter 3 in [6], as shown in Theorem 4.
Theorem 4
@@
$\begin{cases}v(\bm{x})<1&\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathcal{X}_{r}\cup(\widehat{\mathcal{X}}\setminus\mathcal{X})\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$ (11)
@@
$\begin{cases}v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha>1\wedge\frac{\beta\alpha}{\alpha-1}>1\\ 1-(1-v(\bm{x}_{0}))(1-\beta)^{N}&\text{if~{}}\alpha>1\wedge\frac{\beta\alpha}{\alpha-1}\leq 1\end{cases}.$ (12)
Proof:
Since $v(\bm{x})\geq 1$ for $\bm{x}\in\mathcal{X}_{r}\cup(\widehat{\mathcal{X}}\setminus\mathcal{X})$, we have
$\begin{split}&\mathbb{P}^{\infty}\left(\begin{aligned} &\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\mathcal{X}_{r}\wedge\\ &\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(l)\in\mathcal{X}\end{aligned}\middle|\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\right)\leq\mathbb{P}^{\infty}(\exists k\in\mathbb{N}_{\leq N}.v(\bm{\phi}^{\bm{x}_{0}}_{\pi}(k))\geq 1\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}).\end{split}$
@@
Also, since $\bm{x}_{0}\in\{\bm{x}\in\widehat{\mathcal{X}}\mid v(\bm{x})<1\}$, Theorem 3 in Chapter 3 in [6] implies that the bounds in (12) are upper bounds for $\mathbb{P}^{\infty}(\exists k\in\mathbb{N}_{\leq N}.v(\bm{\phi}^{\bm{x}_{0}}_{\pi}(k))\geq 1\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r})$ and thus the conclusion holds. ∎
Following the comparison between constraints (3) and (5), we can conclude that constraint (9) is more expressive that constraint (11).
Remark 4
@@
Inspired by Theorem 3 in Chapter 3 in [6], $c$-martingale for stochastic discrete-time systems was proposed for temporal verification in [5]. Under the assumption that $\bm{f}(\cdot,\cdot)\colon\mathcal{X}\times\Theta\rightarrow\mathcal{X}$ (i.e., $\mathcal{X}$ is a robust invariant set for system (1)), the $c$-martingale in [5] is to certify upper bounds of the probability with which system (1) starting from $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$ will reach the set $\mathcal{X}_{r}$ within a specified bounded time horizon. Under this strong assumption, $c$-martingale satisfies (9) in Theorem 3 with $\alpha=1$. This assumption is also imposed in [17]. It is worth noting here that $\widehat{\mathcal{X}}$ in our conditions is not needed if the assumption that $\bm{f}(\cdot,\cdot)\colon\mathcal{X}\times\Theta\rightarrow\mathcal{X}$ holds.
IV-B Sufficient Conditions for Lower Bounds

### v2

@@ 4 Finite-time Reach-avoid Verification
Definition 4.
[…]
@@
    $\widehat{\mathcal{L}}=\{1,2,3\}$ is a set of three locations;
@@
    with $\widehat{\bm{f}}_{1}(\bm{x},\bm{\theta})=\bm{f}(\bm{x},\bm{\theta})$, $\widehat{\bm{f}}_{2}(\bm{x},\bm{\theta})=\bm{x}$ and $\widehat{\bm{f}}_{3}(\bm{x},\bm{\theta})=\bm{x}$, and $1_{\widehat{\mathcal{X}}_{i}}(\bm{x})$ is the indicator function of the set $\widehat{\mathcal{X}}_{i}$, i.e., $1_{\widehat{\mathcal{X}}_{i}}(\bm{x})=1$ if $\bm{x}\in\widehat{\mathcal{X}}_{i}$; otherwise, $1_{\widehat{\mathcal{X}}_{i}}(\bm{x})=0$,
@@
    $\widehat{\Omega}=\{\bm{x}^{\prime}\in\mathbb{R}^{n}\mid\bm{x}^{\prime}=\bm{f}(\bm{x},\bm{\theta}),\bm{x}\in\mathcal{X},\bm{\theta}\in\Theta\}\cup\mathcal{X};$
@@
[…]
The trajectory of system (8), induced by initial state $\bm{x}_{0}\in\widehat{\mathcal{X}}$ and disturbance policy $\pi$, is denoted by $\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$.
@@
4.1 Sufficient Conditions for Upper Bounds
@@
The sufficient condition requires the existence of a non-negative function $v(\bm{x})$ defined on the set $\widehat{\mathcal{X}}$ that satisfies two key properties: (1) it must be greater than or equal to one on the target set $\mathcal{X}_{r}$, and (2) its $\alpha$-scaled expected value at the next step under the dynamics in (1) should increase by at most $\alpha\beta$ relative to its current value for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$, where $\alpha\in(0,1]$, and $\beta\in[0,1]$. Upper bounds are derived as follows. Lemma 2 equates the probability of system (8) entering $\mathcal{X}_{r}$ at $t=N$ with the cumulative probability of system (1) reaching $\mathcal{X}_{r}$ by time $N$ without leaving $\mathcal{X}$ before hitting the target. Then, we derive upper bounds using system (8) by analyzing $\alpha$. Since system (8) remains stationary when starting from $\widehat{\mathcal{X}}\setminus\mathcal{X}$ and $\mathcal{X}_{r}$, implying $\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]=v(\bm{x})$ for $\bm{x}\in(\widehat{\mathcal{X}}\setminus\mathcal{X})\cup\mathcal{X}_{r}$, we reformulate the constraint on $v(\bm{x})$ using system (8): its $\alpha$-scaled expected value at the next step along (8) grows by at most $\alpha\beta$ relative to its current value across the entire set $\widehat{\mathcal{X}}$. For $\alpha\in(0,1)$ and $\alpha=1$, the upper bounds follow from recursively applying this reformulated constraint.
Theorem 3.
@@
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta,&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r},\\ v(\bm{x})\geq 1,&\forall\bm{x}\in\mathcal{X}_{r},\\ v(\bm{x})\geq 0,&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}_{r},\end{cases}$ (9)
@@
$\begin{cases}v(\bm{x}_{0})+\beta N,&\text{if~}\alpha=1\wedge\beta\in[0,1],\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1},&\text{if~}\alpha\in(0,1)\wedge\beta\in[0,1].\end{cases}$
Proof.
@@
[…]
@@
Thus,
$\begin{split}&\mathbb{P}^{\infty}\big(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big)=\mathbb{E}^{\infty}[1_{\mathcal{X}_{r}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))]\\ &\leq\alpha^{-N}v(\bm{x}_{0})+\alpha\beta\frac{(1-\alpha^{-N})}{\alpha-1}.\end{split}$
According to Lemma 2, we have the conclusion.
@@
On the other hand, the analysis for the upper bounds in Theorem 3 is similar to the one in Theorem 1. When $\alpha=1$ and $\beta=0$, the upper bound is $v(\bm{x}_{0})$. It is an upper bound of the probability $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big)$, which is the probability of reaching the target set $\mathcal{X}_{r}$ eventually while staying within the safe set $\mathcal{X}$ before the first target hitting time.
Remark 3.
In Theorem 3, $\alpha$ is limited to be in $(0,1]$. We can obtain upper bounds for the case with $\alpha>1$ according to Theorem 3 in Chapter 3 in [17] or Proposition 1 in Section 3, as shown in Corollary 1.
Corollary 1.
@@
[…]
@@
$\begin{cases}v(\bm{x}_{0})+\beta N,&\text{if~}\alpha=1,\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1},&\text{if~}\alpha>1\wedge\frac{\beta\alpha}{\alpha-1}>1,\\ 1-(1-v(\bm{x}_{0}))(1-\beta)^{N},&\text{if~}\alpha>1\wedge\frac{\beta\alpha}{\alpha-1}\leq 1.\end{cases}$ (12)
Proof.
Like in [26], consider the stopped version of the stochastic process satisfying (1), which ceases evolving upon exiting the set $\mathcal{X}$. Given a disturbance signal $\pi$, the trajectory of the stopped process, starting from $\bm{x}_{0}$, is denoted by $\{\bar{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i)\}_{i\in\mathbb{N}}$, which satisfies
$\begin{cases}\bar{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i+1)=1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\bar{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i))+1_{\mathcal{X}}(\bar{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i))\bm{f}(\bar{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i),\bm{\theta}(i)),\\ \bar{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(0)=\bm{x}_{0}.\end{cases}$
It is easy to conclude that any sample trajectory for the stopped process starting from any state in $\widehat{\mathcal{X}}$ cannot leave the set $\widehat{\mathcal{X}}$, and $\{\pi\mid\exists k\in\mathbb{N}_{\leq N}.\bar{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\}$$\{\pi\mid\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\bigwedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\}$ for $\bm{x}_{0}\in\widehat{\mathcal{X}}\setminus\mathcal{X}$. Therefore, the probability $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}_{\leq N}.\bar{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big)$ of reaching $\mathcal{X}_{r}$ for the stopped process is equivalent to the reach-avoid probability $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\bigwedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big)$.
According to Theorem 3 in Chapter 3 in [17] (or, Proposition 1 in Section 3), we have that if there exist a continuous non-negative function $v(\bm{x})\colon\widehat{\mathcal{X}}\rightarrow\mathbb{R}$, $\alpha\in[1,\infty)$, and $\beta\in[0,1]$ such that
[…]
@@
then $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}_{\leq N}.\bar{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\big)$ has upper bounds in (12), and thus $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\bigwedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big)$ has upper bounds in (12).
On the other hand, since $\bar{\bm{\phi}}_{\pi}^{\bm{x}}(1)=\bm{x}$, $\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X},\forall\pi$, and $\bar{\bm{\phi}}_{\pi}^{\bm{x}}(1)=\bm{\phi}_{\pi}^{\bm{x}}(1)$, $\forall\bm{x}\in\mathcal{X},\forall\pi$, we obtain that constraint (13) is equivalent to (11). Thus, the conclusion holds and the proof is completed. ∎
Following the comparison between constraints (3) and (5), we conclude that if there exist a function $v(\bm{x})$, $\alpha\geq 1$, and $\beta\in[0,1]$ satisfying (11), then the same function $v(\bm{x})$ with $\frac{1}{\alpha}$ and $\beta\in[0,1]$ will satisfy (9). Additionally, it is noted that a similar constraint, associated with the same upper bounds in (12), has been presented in Proposition 2 of [26], which is also derived from Theorem 3 in Chapter 3 of [17] (or, Proposition 1). [26] utilized this constraint to determine upper bounds of the probability of reaching the set $\mathcal{X}_{r}$ for stopped processes that cease evolving upon exiting the interior of the set $\mathcal{X}$. The stopped process is the same as the one in the proof of Corollary 1 when the set $\mathcal{X}$ is open. As shown in the proof of Corollary 1, the probability of reaching $\mathcal{X}_{r}$ for the stopped process in [26] is equivalent to the reach-avoid probability $\mathbb{P}^{\infty}\big(\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\bigwedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big)$ when the set $\mathcal{X}$ is open. This indicates that the safety verification problem in [26] for stopped processes is actually a reach-avoid verification problem for the original process in the present work. Although the constraint in [26] and the one (11) are constructed based on the same stopped process (when the set $\mathcal{X}$ is open) and Theorem 3 in Chapter 3 of [17], constraint (11) is more stringent than the one in [26], including additional conditions of $(1-\frac{1}{\alpha})v(\bm{x})-\beta\leq 0,\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}$ and $v(\bm{x})\geq 0,\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}$. To the best of my knowledge, I believe that these additional constraints are necessary and cannot be omitted according to Theorem 3 in Chapter 3 of [17] (or, Proposition 1).
Remark 4.
@@
Under the assumption that $\bm{f}(\cdot,\cdot)\colon\mathcal{X}\times\Theta\rightarrow\mathcal{X}$ (i.e., $\mathcal{X}$ is a robust invariant set  [^1] Generally, a robust invariant set is fundamentally challenging, even impossible to compute even if it exists. for system (1)), the $c$-martingale was employed in [13] to certify upper bounds of the probability with which system (1) starting from $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$ will reach the set $\mathcal{X}_{r}$ within a specified bounded time horizon. Under this strong assumption, $c$-martingale satisfies (9) in Theorem 3 with $\alpha=1$. This assumption is also imposed in [38]. It is worth noting that $\widehat{\mathcal{X}}$ in our conditions is not needed if $\bm{f}(\cdot,\cdot)\colon\mathcal{X}\times\Theta\rightarrow\mathcal{X}$ holds.
4.2 Sufficient Conditions for Lower Bounds